Question 3

Run the following code and explain, using your own words, what is happening with each

variable (w,x,y,z) and what is the difference between ++x and w++ (is there any

influence on the result between prefix or postfix ++ operator?). Provide a snippet of the

output showing the result.

int main()

{

int w = 20, x = 20;

int y = 5, z = 5;

y = y + (++x);

z = z + (w++);

cout << "x=" << x << " and y=" << y << endl;

cout << "w=" << w << " and z=" << z << endl;

system("pause");

return 0;

}

Provide a paragraph that answers question 3.

Provide 4 valid test cases for (w, x, y, z).

Test #

Valid / Invalid Data

Description of test

Input Value

Actual Output

Test Pass / Fail

1

Valid

W

Pass

2

Valid

X

Pass

3

Valid

Y

Pass

4

Valid

Z

Pass
Answer #1

The values of the above program after execution is

x = 21

y = 26

w = 21

z = 25

The difference between ++x and w++ is

For ++x the value is incremented first and assigns to the variable x..That's why y = y + (++x) here y = 26 because first y is equal to 5 plus x value is 21(pre increment).So 5 + 21 =26.

For w++ the value for w is assigned first and increments the value.That's why z = z + (w++); here z = 25 because the first z value is equal to 5 plus w value is 20(post increment).So 5 +20 = 25

Now the z value becomes incremented so the z value is 21.

The x value is also updated after incremented so the value of x is 21